Circular No. 7586

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GRB 010222 K. Z. Stanek, Harvard-Smithsonian Center for Astrophysics (CfA); P.

Garnavich, University of Notre Dame; and S. Jha and M. Pahre, CfA, report: “CCD imaging of the error box for the bright γ-ray burst GRB 010222 reported by BeppoSax (GCN 960; available at http://gcn.gsfc. nasa.gov/gcn/gcn3/960.gcn3) confirms the optical afterglow detected by Henden (GCN 962). Images were obtained by J. McDowell and R. Kilgard (CfA) using the Fred L. Whipple Observatory (FLWO) 1.2-m telescope beginning Feb. 22.46 UT and show the transient at R = 18.2 ± 0.1. The object faded by 0.18 ± 0.04 mag over 80 min of observation. The position (determined from HST Guide Star Catalog star positions) of the optical candidate for GRB 010222 is α = 14^{h}52^{m}12^{s}.55, δ = +43^{o}01′06′′.3 (equinox 2000.0), with an uncertainty of 0′′.3. Spectra (range 362–756 nm) of the afterglow were obtained by M. Calkins (FLWO) with the 1.5-m Tillinghast telescope (+ FAST spectrograph), starting at Feb. 22.51. The data show a blue continuum with many narrow absorption features clearly present. Preliminary identification of the strongest features as Mg II and Fe II suggests the presence of at least one absorption system at z = 1.476. The absorption may be from interstellar gas in the host galaxy, but its presence at least places a lower limit on the redshift of the GRB.”

COMET C/2000 W1 (UTSUNOMIYA-JONES) C. W. Hergenrother, Lunar and Planetary Laboratory, reports that

this comet has undergone a rapid fading, with R-band photometry showing m_{1} ∼ 16.5 for a 1′.7 coma on a co-added 2400-s CCD exposure taken on Feb. 12.6 UT with the Catalina 1.5-m reflector. No nuclear condensation was visible to a limiting mag of 21.0. Earlier visual m_{1} estimates: Jan. 17.86, 10.1 (Y. Nagai, Yamanashi, Japan, 0.32-m reflector); 22.88, 10.5 (K. Yoshimoto, Yamaguchi, Japan, 0.25-m reflector); 28.77, 12.0: (M. Mattiazzo, Wallaroo, S. Australia, 0.20-m reflector); 30.28, 11.6 (P. M. Raymundo, northwest of Salvador, Brazil, 0.25-m reflector).

V445 PUPPIS Visual magnitude estimates: Jan. 29.386 UT, 9.6 (J. Bedient, Hon-

olulu, Hawaii); Feb. 1.787, 9.8 (A. Pearce, Nedlands, W. Australia); 7.777, 10.1 (Pearce); 14.116, 9.7 (S. Otero, Buenos Aires, Argentina); 18.06, 10.0 (J. Carvajal, Avila, Spain).

2001 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green